Consider the configuration of three point charges (all having the same charge) forming a equilateral triangle (all sides have length a) in the x-y plane. Assume that all three charges are released from rest at the same time. When the charges are far away from each other. What is their total kinetic energy ?


1
Expert's answer
2021-08-16T08:34:14-0400
"K=E_p\\K=3k\frac{q^2}{d}\\K=3\sqrt{3}k\frac{q^2}{a}"